OpenStudy (anonymous):

Explain why the number 7^20+7^21+7^22 could not be prime

OpenStudy (amistre64):

cause its chock full of multiplication for starters :)

OpenStudy (amistre64):

but thats prolly not why

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

At first I thought it was some great number theory with mods/fermat's little theorem etc, but it isn't:/ 7^20+7^21+7^22 = 7(7^19 + 7^20 + 7^22) = ab where a,b > 1 so is not prime

OpenStudy (anonymous):

Thank You . ( :

OpenStudy (amistre64):

i knew it..... there was a multiplication involved lol